OpenStudy (anonymous):

Solve the proportion. 5/x - 3 = 10/x

OpenStudy (anonymous):

@dc104

OpenStudy (anonymous):

(5/x - 3)x = 10 5 - 3x = 10 3x = -5 x = -5/3

OpenStudy (anonymous):

Back to how we were doing it before. We don't have to move any constants over. This time, we only have to deal with the multiplication (5/x -3)x, which is easy enough. the 5/x*x cancel to leave 5 and of course 3*x = 3x.